please i need help with this problem find the maximum value of y= -x^2+6x
Answer by stanbon(75887)

find the maximum value of y= -x^2+6x
The equation is a quadratic with a=-1,b=6
The maximum or minimum point of the graph is at x=-b/2a
-b/2a = -6/-2 = 3
The y value when x=3 is -3^2+6*-3= -9-18=-27
So the maximum point is at (-3,-27)
=========
Cheers,
stan H.